### v1

@@ 1.3 Our main result
(1.3) $\widetilde{C}=\mathrm{dim}(V_{\rho})\chi(X)(2\zeta^{\prime}(-1)-\log\sqrt{2\pi})+\sum_{j=1}^{r}\frac{\log\nu_{j}}{2\nu_{j}}\sum_{k=1}^{\nu_{j}-1}\frac{\mathrm{Tr}(\rho(\gamma_{j})^{k})e^{i\pi km/\nu_{j}}}{\sin^{2}\left(\frac{\pi k}{\nu_{j}}\right)}$
@@
$\lim_{N\to\infty}\frac{\widetilde{C}}{2N}=\chi(X)(2\zeta^{\prime}(-1)-\log\sqrt{2\pi}).$
@@ 2.2 The relation between $\pi_{1}(X_{1})$ and $\pi_{1}(X)$
Let $\rho:\pi_{1}(X_{1})\to\operatorname{GL}(V_{\rho})$ be a finite-dimensional complex representation of $\pi_{1}(X_{1})$. As the notation suggests, it is not assumed that $\rho$ is unitary. Since $u\in\pi_{1}(X_{1})$ is central, $\rho(u)$ commutes with $\rho(\gamma)$ for all $\gamma\in\pi_{1}(X_{1})$. If $\rho$ is irreducible, $\rho(u)$ must be a scalar multiple of the identity by Schur’s Lemma, say $\rho(u)=\lambda\,I_{n}$. Let $N=\operatorname{lcm}(1,\nu_{1},\ldots,\nu_{r})$ be the least common multiple of $1$ and the orders of the elliptic conjugacy classes. By definition, the orbifold Euler characteristic $\chi(X)$ of $X$ is given by
@@ 3.2 The twisted Bochner–Laplacian
where $\nabla^{\operatorname{LC}}$ is the Levi–Civita connection on $TX$. The negative of the trace of the second covariant derivative is the corresponding Bochner–Laplacian, which we write as
@@ 6 Expressing the determinant in terms of the twisted Selberg zeta function
In order to compute $I^{+}(s,m)$, consider the integral of $F^{+}(z)$ along the contour $C^{+}(R)$ which consists of the real line from 0 to $R$, then along the circle $|z|=R$ with $\arg z$ ranging from 0 to $\pi/2$, and then along the imaginary line segment $\arg z=\pi/2$ from $iR$ to 0. Actually, this is a regularized integral since the function
@@
where $\delta_{m=1}$ equals $1$ if $m=1$ and equals zero, otherwise.
@@ 7 Analysis of the torsion factor for some representations
for $k=1,\ldots,\nu_{j}-1$. If $\nu_{j}\mid N$ then (7.1) is zero, so let us now assume $\nu_{j}$ is not a factor of $N$. Let $r_{j}(N):=2N-\nu_{j}\lfloor\frac{2N}{\nu_{j}}\rfloor$ where $\lfloor x\rfloor$ denotes the integer part of a real number $x$. Moreover, set
@@
(7.2) $\widetilde{C}=2N\chi(X)\left(2\zeta^{\prime}(-1)-\log\sqrt{2\pi}\right)+\sum_{j=1}^{r}\frac{\log\nu_{j}}{2\nu_{j}}\sum_{r_{j}\in A_{j}(N)}\frac{\nu_{j}^{2}-6\nu_{j}r_{j}+6r_{j}^{2}-1}{3}.$
@@
(7.3) $\lim_{N\to\infty}\frac{\widetilde{C}}{2N}=\chi(X)\left(2\zeta^{\prime}(-1)-\log\sqrt{2\pi}\right),$
@@
With the notation as above consider a sequence of representations $\{\rho_{N}\}_{N\geq 1}$, such that $\mathrm{Tr}(\rho_{N}(\gamma_{j}))$ is uniformly bounded by some constant $C(X)$ which is independent of $j$ and $N$.Assume also that the sequence of dimensions $\{\dim(V_{\rho_{N}})\}_{N\geq 1}$ of representations $\rho_{N}$ is such that $\lim_{N\to\infty}\dim(V_{\rho_{N}})=\infty$. We then have that
$\lim_{N\to\infty}\frac{\widetilde{C}}{\dim(V_{\rho_{N}})}=\chi(X)\left(2\zeta^{\prime}(-1)-\log\sqrt{2\pi}\right).$
@@ 8.2 Determinant of the twisted Laplacian associated to a cyclic representation
(8.6) $\widetilde{C}_{0}=\dim(V_{\rho})\chi(X)(2\zeta^{\prime}(-1)-\log\sqrt{2\pi})+\sum_{j=1}^{r}\frac{\log\nu_{j}}{2\nu_{j}}\sum_{k=1}^{\nu_{j}-1}\frac{\mathrm{Tr}(\rho(\gamma_{j})^{k})}{\sin^{2}\left(\frac{\pi k}{\nu_{j}}\right)}.$

### v2

@@ 1.3 Our main result
(1.3) $\widetilde{C}=-\mathrm{dim}(V_{\rho})\chi(X)(2\zeta^{\prime}(-1)-\log\sqrt{2\pi})+\sum_{j=1}^{r}\frac{\log\nu_{j}}{2\nu_{j}}\sum_{k=1}^{\nu_{j}-1}\frac{\mathrm{Tr}(\rho(\gamma_{j})^{k})e^{i\pi km/\nu_{j}}}{\sin^{2}\left(\frac{\pi k}{\nu_{j}}\right)}$
@@
$\lim_{N\to\infty}\frac{\widetilde{C}}{2N}=-\chi(X)(2\zeta^{\prime}(-1)-\log\sqrt{2\pi}).$
@@ 2.2 The relation between $\pi_{1}(X_{1})$ and $\pi_{1}(X)$
Let $\rho:\pi_{1}(X_{1})\to\operatorname{GL}(V_{\rho})$ be a finite-dimensional complex representation of $\pi_{1}(X_{1})$. As the notation suggests, it is not assumed that $\rho$ is unitary. Since $u\in\pi_{1}(X_{1})$ is central, $\rho(u)$ commutes with $\rho(\gamma)$ for all $\gamma\in\pi_{1}(X_{1})$. If $\rho$ is irreducible, $\rho(u)$ must be a scalar multiple of the identity by Schur’s Lemma, say $\rho(u)=\lambda\,I_{n}$. Let $N=\operatorname{lcm}(1,\nu_{1},\ldots,\nu_{r})$ be the least common multiple of 11 and the orders of the elliptic conjugacy classes. By definition, the orbifold Euler characteristic $\chi(X)$ of $X$ is given by
@@ 3.2 The twisted Bochner–Laplacian
where $\nabla^{\operatorname{LC}}$ is the Levi–Civita connection on TXTX. The negative of the trace of the second covariant derivative is the corresponding Bochner–Laplacian, which we write as
@@ 6 Expressing the determinant in terms of the twisted Selberg zeta function
In order to compute $I^{+}(s,m)$, consider the integral of $F^{+}(z)$ along the contour $C^{+}(R)$ which consists of the real line from 0 to $R$, then along the circle $|z|=R$ with $\arg z$ ranging from 0 to $\pi/2$, and then along the imaginary line segment $\arg z=\pi/2$ from iRiR to 0. Actually, this is a regularized integral since the function
@@
where $\delta_{m=1}$ equals 11 if $m=1$ and equals zero, otherwise.
@@ 7 Analysis of the torsion factor for some representations
for $k=1,\ldots,\nu_{j}-1$. If $\nu_{j}\mid N$ then (7.1) is zero, so let us now assume $\nu_{j}$ is not a factor of NN. Let $r_{j}(N):=2N-\nu_{j}\lfloor\frac{2N}{\nu_{j}}\rfloor$ where $\lfloor x\rfloor$ denotes the integer part of a real number $x$. Moreover, set
@@
(7.2) $\widetilde{C}=-2N\chi(X)\left(2\zeta^{\prime}(-1)-\log\sqrt{2\pi}\right)+\sum_{j=1}^{r}\frac{\log\nu_{j}}{2\nu_{j}}\sum_{r_{j}\in A_{j}(N)}\frac{\nu_{j}^{2}-6\nu_{j}r_{j}+6r_{j}^{2}-1}{3}.$
@@
(7.3) $\lim_{N\to\infty}\frac{\widetilde{C}}{2N}=-\chi(X)\left(2\zeta^{\prime}(-1)-\log\sqrt{2\pi}\right),$
@@
With the notation as above consider a sequence of representations $\{\rho_{N}\}_{N\geq 1}$, such that $\mathrm{Tr}(\rho_{N}(\gamma_{j}))$ is uniformly bounded by some constant $C(X)$ which is independent of $j$ and NN.Assume also that the sequence of dimensions $\{\dim(V_{\rho_{N}})\}_{N\geq 1}$ of representations $\rho_{N}$ is such that $\lim_{N\to\infty}\dim(V_{\rho_{N}})=\infty$. We then have that
$\lim_{N\to\infty}\frac{\widetilde{C}}{\dim(V_{\rho_{N}})}=-\chi(X)\left(2\zeta^{\prime}(-1)-\log\sqrt{2\pi}\right).$
@@ 8.2 Determinant of the twisted Laplacian associated to a cyclic representation
(8.6) $\widetilde{C}_{0}=-\dim(V_{\rho})\chi(X)(2\zeta^{\prime}(-1)-\log\sqrt{2\pi})+\sum_{j=1}^{r}\frac{\log\nu_{j}}{2\nu_{j}}\sum_{k=1}^{\nu_{j}-1}\frac{\mathrm{Tr}(\rho(\gamma_{j})^{k})}{\sin^{2}\left(\frac{\pi k}{\nu_{j}}\right)}.$